### v1

# Establishing a Baseline for Gaze-driven Authentication Performance in VR: A Breadth-First Investigation on a Very Large Dataset

Abstract
This paper performs the crucial work of establishing a baseline for gaze-driven authentication performance to begin answering fundamental research questions using a very large dataset of gaze recordings from $9202$ people with a level of eye tracking (ET) signal quality equivalent to modern consumer-facing virtual reality (VR) platforms. The size of the employed dataset is at least an order-of-magnitude larger than any other dataset from previous related work. Binocular estimates of the optical and visual axes of the eyes and a minimum duration for enrollment and verification are required for our model to achieve a false rejection rate (FRR) of below $3$% at a false acceptance rate (FAR) of $1$ in $50000$. In terms of identification accuracy which decreases with gallery size, we estimate that our model would fall below chance-level accuracy for gallery sizes of $148000$ or more. Our major findings indicate that gaze authentication can be as accurate as required by the FIDO standard when driven by a state-of-the-art machine learning architecture and a sufficiently large training dataset.
1 Introduction
Accurate, spoof-resistant, and privacy-aware user authentication on extended reality (XR) devices is an important topic of research necessary to provide the highest levels of security to users of such devices. Passwords can be difficult to remember and inconvenient to enter, and they are not suitable for continuous authentication. Comparatively, eye gaze authentication [22] might reduce cognitive load and increase convenience. For the best user experience and highest security, gaze signatures would need to exhibit high distinctiveness, permanence, collectability, and universality [19], and they would also need to be highly spoof-resistant and enable fast accept/reject decisions.
Eye gaze authentication can leverage existing eye tracking (ET) hardware in modern XR devices to provide accurate and spoof-resistant unlocking of such devices. It is also important to state that while devices such as the Apple Vision Pro [2] employ fine-tuned hardware for user authentication via iris patterns, iris-based authentication is not always possible for gaze estimation pipelines that employ non-image sensors [1, 4, 14, 38, 41] or for image-based gaze estimation pipelines with optics and algorithms that are fine-tuned to increase ET signal quality and reduce power consumption at the expense of the visibility of fine-grained information related to iris ridges [3, 30]. Another advantage of gaze authentication is its high spoof-resistance [29, 34, 36] due to incorporating not only static parameters of gaze such as instantaneous measurement of a person’s optical and visual axes but also the dynamics of gaze which includes the variety and complexity of various eye movement types [24].
In this work, we report on various aspects of user authentication using the GazePro dataset. GazePro contains ET signals from a population of $9202$ people recorded at $72$ Hz using a virtual reality (VR) headset that provides a level of ET signal quality equivalent to a Meta Quest Pro [6, 40]. Our use of such hardware is unlike many previous studies [27, 29] which often utilize laboratory-grade ET systems with signal quality characteristics that are currently unachievable in consumer-grade XR devices.
We report a best-case result of $2.4$% false rejection rate (FRR) at a $1$-in-$50000$ false acceptance rate (FAR) when using $20$ seconds of gaze data during a random saccade (jumping dot) task. To the best of our knowledge, this is the first time such a high level of authentication accuracy has been achieved via gaze alone. Thus, we are the first to demonstrate that gaze can potentially provide meaningful user authentication performance under FIDO guidelines [37]. Importantly, this result was achieved using the current state-of-the-art machine learning architecture, Eye Know You Too (EKYT) [27], signifying the importance of large-scale datasets for achieving meaningful performance at the level of ET signal quality employed in our work. It is also important to note that identification rates decrease with increasing gallery sizes, and we roughly estimate that we might reach random-chance identification rates starting at a gallery size of about $148000$ identities.
Here we perform a breadth-first investigation of several fundamental research questions to provide necessary baselines for future work in this important research area and to begin answering several of these questions for the first time: (RQ1) Does the use of binocular gaze improve authentication accuracy over monocular gaze? (RQ2) Does the use of optical axis estimates in addition to visual axis estimates improve authentication accuracy? (RQ3) What aspects of training the state-of-the-art machine learning architecture result in the best authentication accuracy? (RQ4) How does the size of the training dataset affect authentication accuracy? (RQ5) How quickly does identification accuracy degrade as the gallery size increases? (RQ6) How does ET signal quality, specifically spatial accuracy, affect authentication accuracy? (RQ7) How does the length of the employed gaze signal affect authentication accuracy? (RQ8) Can authentication accuracy be improved by increasing the duration of the enrollment session? (RQ9) Can gaze authentication be task-independent? (RQ10) Is the permanence of the learned embedding features correlated with authentication accuracy?
2 Background
Since the seminal work of Kasprowski & Ober [22], gaze-based authentication has been studied extensively [5, 7, 9, 10, 13, 18, 20, 21, 23, 25, 27, 29, 35, 43]. End-to-end deep learning approaches have only recently been explored [5, 18, 20, 27, 29], with the current state-of-the-art model being EKYT [27]. The EKYT architecture and training methodology are employed for all experiments in the present study. Some modifications to the original training parameters are also explored in pursuit of even better performance on the employed dataset.
Studies in this domain typically employed either monocular gaze data (such as [27]) or binocular gaze data (such as [29]), but the literature currently lacks studies that compare performance with monocular versus binocular data. Binocular data contains information that cannot be measured from monocular data alone, such as eye dominance and vergence angle. Such additional information might enable increased authentication accuracy using binocular data compared to monocular data. The present study provides such a performance comparison.
The vast majority of studies in this domain employed estimates of the visual axis (sometimes called the line of sight) to represent gaze. With current ET technology which cannot directly measure the position of the fovea, the visual axis can be estimated only after performing a user-specific calibration procedure. But there are scenarios such as continuous authentication in which it is impractical to precede every verification attempt with a calibration procedure. In contrast, the optical axis (sometimes called the pupillary axis) can be estimated without a user-specific calibration procedure [8]. The optical and visual axes are illustrated in Figure 1. Even in scenarios where the visual axis can be estimated, perhaps a combination of both the optical and visual axes might improve authentication accuracy. To the best of our knowledge, the present study is the first to assess gaze-based authentication when using the optical axis and when using a combination of both the optical and visual axes.
Figure 1: A simplified model of the eye, visualizing the imaginary optical and visual axes. The optical axis passes through the eyeball center and the pupil center. The visual axis connects the center of the fovea (“foveola”) and the gaze target (“object of regard”).
Currently, the largest public ET datasets used for gaze authentication contain on the order of hundreds of people (e.g., $150$ in JuDo1000 [28], $322$ in GazeBase [15], and $407$ in GazeBaseVR [26]). In contrast, the present study employs GazePro, a dataset of $9202$ people, which is an order-of-magnitude larger than existing datasets. The GazePro dataset affords the present study the unique opportunity to explore the applicability of reliable gaze authentication on a never-before-seen scale.
There are several characteristics commonly used to describe ET signal quality [17], such as sampling rate, spatial accuracy (error), and spatial precision (jitter). Some studies have investigated how different sampling rates and levels of spatial precision affect authentication performance [27, 31], with the general consensus being that lower sampling rates and/or worse spatial precision lead to worse authentication performance. There is also evidence [32] that high-frequency (“noise”) components of gaze signals carry person-specific information with features that are reliable across short test-retest intervals. It is currently hypothesized that the gaze estimation pipeline creates person-specific noise. But the effect of spatial accuracy on authentication performance is underexplored. A recent study by Raju et al. [33] suggested that authentication performance does vary across different levels of spatial accuracy. Surprisingly, that study reported that they achieved better authentication performance on a subset of users from GazeBaseVR [26] with worse spatial accuracy than on those with better spatial accuracy. The present study further explores the effect of spatial accuracy on authentication performance across a larger user population.
Test-set population sizes vary across studies in this domain, from $25$ [29] to $59$ [27] to $149$ [10]. Few studies meet the FIDO guidelines [37] of a test-set population size of $245$ users. Further, the theoretical analysis by Friedman et al. [12] used synthetic feature distributions to show that verification accuracy remained stable across population sizes while identification accuracy decreased with larger population sizes. Until now, no study has had a large enough dataset size to confirm these findings in practice in the ET domain.
Similarly, the minimum duration of each verification attempt also widely varied across studies, from $1$ second [29] to $5$ seconds [27] to $60$ seconds [10]. As one might expect, studies have found that authentication accuracy generally improved as more data was used for each verification attempt [27, 29]. The study by Lohr & Komogortsev [27] used the same amount of data for both enrollment and verification and found that at least $30$ seconds of gaze data was necessary to achieve their best performance of $5.08$% FRR at a $1$-in-$10000$ FAR. The study by Makowski et al. [29] instead used a fixed enrollment duration of $24$ seconds and only varied the verification duration. The present study further explores how authentication accuracy is affected by data length across a much larger population of people.
Most studies in this domain trained and evaluated on the same task, such as a reading task (e.g., [10]) or a random saccade (jumping dot) task (e.g., [29]). A study by Holland & Komogortsev [16] was one of the first to demonstrate that similar gaze authentication performance could be achieved on different tasks. Lohr & Komogortsev [27] later demonstrated that a model trained across a variety of different tasks could achieve similar authentication performance across all such tasks, including an additional task that was not part of the training set. The present study further explores the task-independence of gaze authentication across a larger user population.
Friedman et al. [11] have previously demonstrated that high temporal persistence (permanence) of learned features is important for improved authentication performance. Raju et al. [33] have also found a high correlation between permanence and authentication performance. Lohr & Komogortsev [27] have shown that gaze signatures can exhibit high permanence over periods as long as $3$ years. The present study provides further insight into the relationship between permanence (on a short test-retest interval) and authentication performance on a dataset with a much larger user population.
3 Methodology
An overview of our methodology is presented here and summarized by the diagram in Figure 2. More details can be found in the supplementary material for this manuscript.
Figure 2: Overview diagram of the methodology for computing similarity scores.
3.1 Dataset
We employed GazePro, an internal dataset of gaze signals collected from $9202$ participants which is an order-of-magnitude more participants than have been used for studies in this domain (e.g., [15, 26, 28]). Having access to such a large population gives us the unique opportunity to evaluate gaze authentication at a previously unattainable scale. The dataset is split into a training set of $6747$ users and a testing set of $2455$ users.
Of the $9202$ participants in GazePro, $3673$ self-reported as male, $5376$ as female, and $153$ as neither male nor female. The participants’ ages ranged from $13$ to $88$ years (median=$33$, IQR=$18$). All participants had normal or corrected-to-normal vision. Data collection went through the appropriate IRB approval protocol and received appropriate consent from participants. Additional participant details can be found in the supplementary material.
Gaze data in GazePro were collected at $72$ Hz on a VR headset with ET signal quality and other general characteristics equivalent to a Meta Quest Pro [6, 40]. Each recording contains binocular estimates of the optical and visual axes. The visual axis is a linear transformation of the optical axis. The parameters of the linear model (essentially, the angle kappa between the optical and visual axes) are fit on a per-user basis via a short calibration procedure.
3.2 Model training
We closely followed the same training methodology that was used for EKYT [27] with only a couple of differences. Models were re-implemented using PyTorch and trained on an NVIDIA V100 GPU with 16 GB VRAM.
[…]
The model is trained using multi-similarity (MS) loss [39], which encourages the model to learn a well-clustered embedding space with smaller distances between embeddings from the same user and larger distances between embeddings from different users. We note that Lohr & Komogortsev [27] included an additional classification loss term, but we exclude that term along with the associated classification layer. This was done because the original work showed that training with MS loss alone slightly improved authentication performance in some instances.
Although the original approach by Lohr & Komogortsev [27] trained $4$ models using a $4$-fold split of the training set, in most of our experiments we instead train a single model on the full training set. We found that training a single model (still using an embedding size of $128$) generally performs better on our testing set than the original approach that concatenates the embeddings from $4$ models.
3.3 Model evaluation
We consider two main evaluation scenarios. In both scenarios, the embedding of a gaze sequence is obtained by first embedding each non-overlapping segment of $5$ seconds and then taking the mean across all such segment embeddings.
The first scenario is verification, for which we try to answer the question, “Is this user who they claim to be?” A potential user claims to be a specific person (e.g., the owner of a VR headset), their gaze signature is captured, and an accept/reject decision must be made based on the similarity of the gaze signatures captured from the claimed participant at enrollment time (e.g., during account creation) and from the potential user at verification time. We measure performance in this scenario using equal error rate (EER) and FRR at $1$-in-$50000$ FAR (denoted $\text{FRR}_{0.002\%}$).
The second scenario is identification, for which we try to answer the question, “Who is that person specifically?” A user’s gaze signature is captured, the similarity scores are computed between that user’s gaze signature and all gaze signatures from the population of enrolled participants, and the participant with the highest similarity is determined to be the best match. It is possible to reject a user if the highest similarity score is below some predetermined threshold, in which case it is assumed the user is not part of the enrolled population; but in our case, we evaluated identification rates only on users who were part of the enrolled population. We measure performance in this scenario using Rank-1 identification rate (IR).
4 Results
Table 1: Verification performance for a selection of experiments of interest. Lower values indicate better performance. The best results are highlighted in gray. To reduce visual clutter, the symbol $\cdot$” is used to indicate that the value is repeated from the row above. In the “Eye” column, “L” is the left eye and “R” is the right eye. In the “Axis” column, “O” is the optical axis and “V” is the visual axis.
Experiment Eye Axis Train set Test set EER (%) $\text{FRR}_{0.002\%}$ (%)
1 L V Full Full $8.53$ $99.2$
2 L,R $\cdot$ $\cdot$ $\cdot$ $3.73$ $93.1$
3 L,R O,V Full Full $0.08$ $17.7$
[…]
7 L,R O,V N=0250 Full $1.59$ $92.2$
8 $\cdot$ $\cdot$ N=0500 $\cdot$ $0.98$ $84.1$
9 $\cdot$ $\cdot$ N=1000 $\cdot$ $0.66$ $82.6$
10 $\cdot$ $\cdot$ N=2000 $\cdot$ $0.61$ $78.9$
11 $\cdot$ $\cdot$ N=3000 $\cdot$ $0.10$ $22.0$
12 $\cdot$ $\cdot$ N=4000 $\cdot$ $0.42$ $71.7$
13 $\cdot$ $\cdot$ N=5000 $\cdot$ $0.07$ $13.7$
14 $\cdot$ $\cdot$ N=6000 $\cdot$ $0.06$ $12.6$
15@$250$ L,R O,V $4$ folds, N=4000 N=0250 $0.42$ $65.9$
15@$500$cdot $\cdot$ $\cdot$ N=0500 $0.45$ $68.9$
15@$1$K $\cdot$ $\cdot$ $\cdot$ N=1000 $0.41$ $70.4$
15@$2$K $\cdot$ $\cdot$ $\cdot$ N=2000 $0.43$ $69.9$
15@$3$K $\cdot$ $\cdot$ $\cdot$ N=3000 $0.44$ $71.3$
15@$4$K $\cdot$ $\cdot$ $\cdot$ N=4000 $0.44$ $71.0$
15@$5$K $\cdot$ $\cdot$ $\cdot$ N=5000 $0.44$ $70.9$
16@HighErr L,R O,V High err. High err. $1.00$ $75.9$
16@MidErr $\cdot$ $\cdot$ $\cdot$ Mid err. $0.41$ $75.9$
16@LowErr $\cdot$ $\cdot$ $\cdot$ Low err. $0.73$ $76.7$
17@HighErr $\cdot$ $\cdot$ Low err. High err. $2.25$ $91.8$
17@MidErr $\cdot$ $\cdot$ $\cdot$ Mid err. $0.41$ $70.6$
17@LowErr $\cdot$ $\cdot$ $\cdot$ Low err. $0.42$ $92.2$
5@5s L,R O,V Full, $759$ epochs, $m=1024$ Full $0.18$ $41.9$
5@10s $\cdot$ $\cdot$ $\cdot$ $\cdot$ $0.08$ $20.6$
5@15s $\cdot$ $\cdot$ $\cdot$ $\cdot$ $0.05$ $8.8$
18@SAC L,R O,V $4$ folds Full $0.10$ $22.9$
18@PUR $\cdot$ $\cdot$ $\cdot$ $\cdot$ $0.10$ $24.8$
We highlight the verification performance under various conditions in Table 1. Unless otherwise noted, results were computed during a random saccade (jumping dot) task with the full training set ($6747$ users) and the full testing set ($2455$ users), using a model trained for $100$ epochs with a minibatch size of $m=256$ samples, and using $20$ seconds of gaze data for both enrollment and verification. In this case, we have $2359$ genuine matches and $5707331$ impostor matches (note that not every user was recorded for one or both of the tasks of interest). We assigned each experiment a unique ID to enable easier reference. Additional experiments can be found in the supplementary material.
Baseline.
Lohr & Komogortsev [27] used measures of the left eye’s visual axis. We did the same in Experiment 1 and achieved an EER of $8.53$% and a $\text{FRR}_{0.002\%}$ of $99.2$%.
(RQ1) Monocular gaze vs binocular gaze.
Although our baseline experiment used data from a single eye, many ET-capable XR systems, including the one employed for GazePro, track both eyes. Additional information, such as eye dominance and vergence angle, is encoded in binocular data. For that reason, authentication accuracy might be increased due to this additional information.
Experiment 2, which used the visual axis from both the left and right eyes, can be compared against our monocular baseline (Experiment 1). This experiment achieved an EER of $3.73$% and a $\text{FRR}_{0.002\%}$ of $93.1$%. These results were substantially better than the monocular baseline, indicating that having access to gaze estimates from both eyes indeed improved authentication performance over having gaze estimates from a single eye.
(RQ2) Optical axis and visual axis.
The majority of studies in this domain use estimates of the visual axis. Working under the condition that we have access to a user’s calibration parameters to estimate the visual axis, what if both the optical and visual axes were fed into the model? This would implicitly provide information to the model about the user’s calibration parameters, so we would expect better performance using both axes than visual axis alone.
Experiment 3, which used both the optical and visual axes from both the left and right eyes, can be compared against our visual axis baseline (Experiment 2). This experiment achieved an EER of $0.08$% and a $\text{FRR}_{0.002\%}$ of $17.7$%. These results were substantially better than the visual axis baseline, which makes sense considering that the user-specific calibration parameters can effectively be inferred from the difference between the two axes.
To further demonstrate the effectiveness of the user-specific calibration parameters, Experiment 4 uses the difference of the two axes (visual minus optical) for both the left and right eyes. This experiment achieves an EER of $0.27$% and a $\text{FRR}_{0.002\%}$ of $45.6$%. These results were still substantially better than the visual axis baseline but worse than when using both axes together.
(RQ3) Training epochs and minibatch size.
We found that increasing the number of training epochs from $100$ to $1000$ yielded a large improvement in performance; however, given the large size of the GazePro dataset, this required more than $3$ weeks to train on a single NVIDIA V100 GPU. Simultaneously increasing the minibatch size from $m=256$ to $m=1024$ further improved performance.
In Experiment 5, we achieved $0.04$% EER and $4.1$% $\text{FRR}_{0.002\%}$ when training for $759$ epochs (terminated early due to time constraints) with a minibatch size of $m=1024$, and when using both the optical and visual axes from both the left and right eyes. These results were substantially better than Experiment 3 which trained for $100$ epochs with a minibatch size of $m=256$. The genuine and impostor similarity score distributions and receiver operating characteristic (ROC) curve for this experiment are presented in Figure 3.
[…]
Our best results were from Experiment 6 where we achieved $0.04$% EER and $2.4$% $\text{FRR}_{0.002\%}$ using $20$ seconds of a random saccade task for both enrollment and verification. To obtain these results, we trained an ensemble of $4$ models using a $4$-fold split of the training set similar to [27]. Each model was trained for $1000$ epochs with a minibatch size of $m=1024$ and used both the optical and visual axes from both the left and right eyes.
(RQ4) Training population size.
Considering that the GazePro dataset has an order-of-magnitude more participants than the largest public ET datasets such as GazeBase [15] and GazeBaseVR [26], we were curious how important it was to have such a large dataset. For these experiments, we trained the model on a smaller subset of subjects from the training set to see how performance changed as the amount of training data increased. The test set is unmodified to facilitate comparisons across experiments.
We observed, perhaps not surprisingly, that performance generally continued to improve as the training population increased from $250$ to $6000$ users (see Figure 4 for a visualization). The improvement is not necessarily monotonic (see for example Experiments 3 and 12), which may be due to the inherent randomness of training and minibatch sampling leading to some models happening to settle on poor local optima. We see the largest improvement when increasing the training population from $2000$ to $3000$ users, but it is difficult to determine whether this was just the first time the model happened to settle on better optima.
Figure 4: Authentication performance vs training population size. Measurements are from Experiments 3 and 7–14.
(RQ5) Testing population size.
An analysis by Friedman et al. [12] found that, using synthetic feature distributions, verification accuracy remained stable as population size increased whereas identification accuracy decreased. Since we have access to real gaze data from thousands of people in the GazePro dataset, we wanted to test whether those results hold in practice.
This experiment (Experiment 15) differs from the others in a few ways. First, we used $4000$ users for the training set and the remaining $5202$ users for the testing set. Second, we trained an ensemble of $4$ models using a $4$-fold split of the training set similar to [27]. Third, we drew $100$ independent samples of $N$ users from the testing set without replacement, ensuring that each selected user had both recordings of interest so they could be both enrolled and verified. We computed performance measures for each sample of $N$ users and then computed the P5 and P95 of each measure across all samples. In Table 1, for the sake of brevity, we reported the average of the P5 and P95 values.
As illustrated in Figure 5, we found that verification rates had a consistent central tendency regardless of population size, whereas the identification rates decreased as the population size increased. These results offer practical validation of the theoretical analysis from Friedman et al. [12]. In Figure 6, we roughly estimated that the identification rate might decrease to below-chance levels when the population size increased to around $148000$ users.
Figure 5: Performance measures vs gallery size from Experiment 15. The shaded region represents the $5$th and $95$th percentiles across $100$ random samples, and the black line is the average of the two percentiles. Top: EER. Bottom: Rank-1 IR.
Figure 6: Identification rate (P95 across $100$ random samples) vs gallery size from Experiment 15. A curve of the form $a\sqrt{x}+b$ is fit to the observations. Extrapolating from this curve, we roughly estimate that the identification rates might reach below-chance levels starting at gallery sizes of around $148000$ users.
(RQ6) Eye tracking signal quality.
ET signal quality over a large population of people varies for many reasons, including but not limited to eye color, eye shape, presence of corrective lenses, makeup, and device placement on a user’s head [17]. Spatial accuracy, which is defined as the error between a user’s actual gaze and the gaze vector estimated by a gaze estimation pipeline, is one of the most important ET signal quality characteristics.
The subset of $8615$ participants who were recorded for both random saccade tasks were partitioned based on their average spatial accuracy during the two tasks. These partitions (illustrated in Figure 7) formed high error and low error train sets each with $3940$ participants, as well as high error, medium error, and low error test sets each with $245$ participants.
Figure 7: Partitions of users in GazePro based on their average spatial accuracy. The degree value above a partition represents the median spatial accuracy across users in that partition. Note that only the subset of $8615$ users who were recorded for both random saccade tasks are included in these partitions.
EER was highest on the high error test set for both Experiments 16 and 17, which may indicate that poor spatial accuracy reduces authentication performance. Additionally, a model trained on data with high/low error performed worse when evaluated on data with low/high error.
(RQ7–8) Enrollment/verification duration.
The experiments described above employed $20$ seconds of data for both enrollment and verification. However, it is important to understand how authentication accuracy changes for shorter presentation durations. In this section, we employed presentation durations of $5$, $10$, and $15$ seconds to establish corresponding performance baselines. To do this, instead of computing a centroid embedding over the first four $5$-second chunks of a recording, we used the first one, two, or three chunks. We used the model from Experiment 5 for this analysis.
Using the first $5$ seconds for enrollment and verification, we achieved $0.18$% EER and $41.9$% $\text{FRR}_{0.002\%}$. Using the first $10$ seconds for enrollment and verification, we achieved $0.08$% EER and $20.6$% $\text{FRR}_{0.002\%}$. Using the first $15$ seconds for enrollment and verification, we achieved $0.05$% EER and $8.8$% $\text{FRR}_{0.002\%}$.
If we instead used a fixed enrollment duration of $30$ seconds and only varied the verification duration, we observed a similar trend of authentication performance improving with increasing verification duration. We observed that enrolling with $30$ seconds and verifying with $X$ seconds (up to $X=15$) yielded roughly similar results as enrolling and verifying with $X+5$ seconds.
(RQ9) Task-independence.
The prior experiments used a random saccade (jumping dot) task for both enrollment and verification. How much, if at all, would performance change by using a smooth pursuit (gliding dot) task instead? We used the same inputs and training/testing sets as Experiment 3 for this analysis, but we trained an ensemble of $4$ models using a $4$-fold split of the training set similar to [27].
With this configuration (Experiment 18), when enrolling and verifying with $20$ seconds from random saccade tasks, we achieved $0.10$% EER and $22.9$% $\text{FRR}_{0.002\%}$. If we instead enrolled and verified with $20$ seconds from smooth pursuit tasks, we achieved $0.10$% EER and $24.8$% $\text{FRR}_{0.002\%}$. These two sets of results were very similar.
(RQ10) Feature permanence.
For this experiment, we employed the model from Experiment 5 to produce embeddings. We then assessed the permanence (temporal persistence, reliability) of L2-normalized embedding features using the intraclass correlation coefficient (ICC), which has been applied to eye movement traits [10, 11, 33] and is also commonly used in other domains (e.g., [42]).
We found that the median ICC of the $128$ embedding features was $0.94$, with a minimum of $0.90$ and a maximum of $0.98$. Even our lowest ICC of $0.90$ meets all but the strictest of the commonly used thresholds [42], so all $128$ embedding features may be considered highly reliable.
5 Discussion
(RQ1) Binocular gaze data yielded a substantial improvement in verification accuracy over monocular gaze data (Experiments 1 and 2). Having access to data from both eyes provides estimates of eye dominance and vergence angle, and the model may be able to utilize this information to increase performance.
(RQ2) We also found that including both the optical and visual axes as inputs to the model yields a substantial improvement in verification accuracy over using the visual axis alone (Experiments 2 and 3). The difference between the visual and optical axes would essentially provide an estimate of a user’s calibration parameters, and this difference alone achieves good verification accuracy (Experiment 4). In a scenario such as continuous authentication where a calibration period could not precede every verification attempt, it may be feasible to employ a cached set of calibration parameters from the enrolled user to estimate the visual axis.
[…]
(RQ4) We observed an increase in verification rates as the training population size increased from $250$ to $6000$ users (Experiments 7–14), with the largest performance improvement occurring between $2000$ and $3000$ users. These results suggest that, if there is a point at which performance stops improving when more users are included in the training set, such a point would be beyond $6000$ users. Also, it is possible that training on an even larger population may yield even better verification rates.
(RQ5) We empirically confirmed in practice that, as the test population size increases, verification rates remain stable whereas identification rates decrease (Experiment 15). Eye movement traits are behavioral rather than physical, so they are expected to lack the distinctiveness necessary for large-scale identification. By fitting a curve of the form $a\sqrt{x}+b$ to the observed decline of identification rates and assuming the trend continues, we might expect random-chance identification rates at gallery sizes of about $148000$ identities (see Figure 6).
(RQ6) Regarding the impact of spatial accuracy on authentication performance, we found that poor spatial accuracy reduced authentication performance (Experiments 16–17), in contrast to a study by Raju et al. [33]. Our results also suggest that, for the best performance, the spatial accuracy of the training dataset should be representative of the level of spatial accuracy present in the testing dataset.
(RQ7–8) We observed that, using the model from Experiment 5, $\text{FRR}_{0.002\%}$ is approximately halved for every additional $5$ seconds used for enrollment and verification, starting at $41.9$% with $5$ seconds and improving to $4.1$% with $20$ seconds. Additionally, increasing the enrollment duration improves performance for shorter verification durations. It is important to note, though, that user calibration parameters have still been determined from a full-length calibration process.
(RQ9) We found that, when using both the optical and visual axes as inputs to the model, enrolling and verifying with a smooth pursuit task results in similar performance as using a random saccade task (Experiment 18). This is evidence that gaze-based verification remains to be task-independent even for larger population sizes.
(RQ10) The learned embedding features of the model from Experiment 5 all have ICCs at or above $0.9$, indicating that all $128$ embedding features are highly reliable.
Further discussion of the limitations of this work can be found in the supplementary material.
6 Conclusion
In this paper, we investigated authentication performance with a large dataset of gaze recordings collected in VR. Our most important findings indicate that the amount of training data is critical to achieve the authentication performance that is expected from modern wearable devices. Given a very large dataset, which in our case was nearly $10000$ people, it is important to have a longer training time and increased minibatch size. We have highlighted that the best authentication performance can be achieved if binocular estimates of both the optical and visual axes are provided. Crucially, and in support of prior theoretical findings, we found that gaze data alone cannot be employed for (re-)identification purposes if the user pool is sufficiently large. We also discovered the impact of spatial accuracy, an important descriptor of ET signal quality, such that higher signal quality resulted in better authentication performance. In addition, we reinforced the importance of data length on resulting accuracy, specifically that longer enrollment/verification durations improved authentication accuracy. We also found that gaze-based authentication in our study appeared to be stimulus- and task-independent. In summary, gaze authentication can be as accurate as required by the FIDO standard when driven by a state-of-the-art machine learning architecture and a sufficiently large training dataset.
Acknowledgments
We thank Meta for providing us with access to the GazePro dataset and compute resources to perform this important work.

### v2

@@
 Data collection went through the appropriate IRB approval protocol and received appropriate consent from participants.
-Additional participant details can be found in the supplementary material.
+Additional details about how the dataset was collected and participant characteristics can be found in the supplementary material.
 Gaze data in GazePro were collected at $72$ Hz on a VR headset with ET signal quality and other general characteristics equivalent to a Meta Quest Pro [6, 40].
-Each recording contains binocular estimates of the optical and visual axes.
+Each recording contains both optical and visual axis estimates from both the left and right eyes.
 The visual axis is a linear transformation of the optical axis.
@@
 In both scenarios, the embedding of a gaze sequence is obtained by first embedding each non-overlapping segment of $5$ seconds and then taking the mean across all such segment embeddings.
-The first scenario is verification, for which we try to answer the question, “Is this user who they claim to be?” A potential user claims to be a specific person (e.g., the owner of a VR headset), their gaze signature is captured, and an accept/reject decision must be made based on the similarity of the gaze signatures captured from the claimed participant at enrollment time (e.g., during account creation) and from the potential user at verification time.
+The first scenario is verification.
+A presenting user claims to be a specific person (e.g., the owner of a VR headset), their gaze signature is captured, and an accept/reject decision must be made based on the similarity of the gaze signatures captured from the person with the claimed identity at enrollment time (e.g., during account creation) and from the presenting user at verification time.
 We measure performance in this scenario using equal error rate (EER) and FRR at $1$-in-$50000$ FAR (denoted $\text{FRR}_{0.002\%}$).
-The second scenario is identification, for which we try to answer the question, “Who is that person specifically?” A user’s gaze signature is captured, the similarity scores are computed between that user’s gaze signature and all gaze signatures from the population of enrolled participants, and the participant with the highest similarity is determined to be the best match.
+The second scenario is identification.
+A user’s gaze signature is captured, the similarity scores are computed between that user’s gaze signature and all gaze signatures from the population of enrolled participants, and the participant with the highest similarity is determined to be the best match.
 It is possible to reject a user if the highest similarity score is below some predetermined threshold, in which case it is assumed the user is not part of the enrolled population; but in our case, we evaluated identification rates only on users who were part of the enrolled population.
@@
 Unless otherwise noted, results were computed during a random saccade (jumping dot) task with the full training set ($6747$ users) and the full testing set ($2455$ users), using a model trained for $100$ epochs with a minibatch size of $m=256$ samples, and using $20$ seconds of gaze data for both enrollment and verification.
-In this case, we have $2359$ genuine matches and $5707331$ impostor matches (note that not every user was recorded for one or both of the tasks of interest).
-We assigned each experiment a unique ID to enable easier reference.
-Additional experiments can be found in the supplementary material.
-Baseline.
-Lohr & Komogortsev [27] used measures of the left eye’s visual axis. We did the same in Experiment 1 and achieved an EER of $8.53$% and a $\text{FRR}_{0.002\%}$ of $99.2$%.
+Not all $2455$ users in the testing set were recorded for one or both of the tasks used for enrollment/verification, so we had fewer than $2455$ genuine matches.
+Instead, in this case we obtained $2359$ genuine matches and $5707331$ impostor matches from a matrix of $2385\times 2394$ match scores.
+We assigned each experiment a unique ID to enable easier reference. Additional experiments can be found in the supplementary material.
+Lohr & Komogortsev [27] used measures of the left eye’s visual axis. As a baseline, we did the same in Experiment 1 and achieved an EER of $8.53$% and a $\text{FRR}_{0.002\%}$ of $99.2$%.
 (RQ1) Monocular gaze vs binocular gaze.
@@
 This experiment achieved an EER of $3.73$% and a $\text{FRR}_{0.002\%}$ of $93.1$%.
-These results were substantially better than the monocular baseline, indicating that having access to gaze estimates from both eyes indeed improved authentication performance over having gaze estimates from a single eye.
+These results were substantially better than the monocular baseline.
 (RQ2) Optical axis and visual axis.
@@
 This experiment achieved an EER of $0.08$% and a $\text{FRR}_{0.002\%}$ of $17.7$%.
-These results were substantially better than the visual axis baseline, which makes sense considering that the user-specific calibration parameters can effectively be inferred from the difference between the two axes.
+These results were substantially better than the visual axis baseline.
 To further demonstrate the effectiveness of the user-specific calibration parameters, Experiment 4 uses the difference of the two axes (visual minus optical) for both the left and right eyes.
@@
 (RQ3) Training epochs and minibatch size.
-We found that increasing the number of training epochs from $100$ to $1000$ yielded a large improvement in performance; however, given the large size of the GazePro dataset, this required more than $3$ weeks to train on a single NVIDIA V100 GPU.
+We found that increasing the number of training epochs from $100$ to $1000$ yielded a large improvement in performance.
 Simultaneously increasing the minibatch size from $m=256$ to $m=1024$ further improved performance.
@@
 These results were substantially better than Experiment 3 which trained for $100$ epochs with a minibatch size of $m=256$.
-The genuine and impostor similarity score distributions and receiver operating characteristic (ROC) curve for this experiment are presented in Figure 3.
-[…]
+The genuine and impostor similarity distributions and receiver operating characteristic (ROC) curve for this experiment are presented in Figure 3.
+Figure 3: Qualitative performance from Experiment 5. The left figure shows histograms of the genuine and impostor similarity score distributions. The right figure shows the ROC curve.
 Our best results were from Experiment 6 where we achieved $0.04$% EER and $2.4$% $\text{FRR}_{0.002\%}$ using $20$ seconds of a random saccade task for both enrollment and verification.
@@
 Spatial accuracy, which is defined as the error between a user’s actual gaze and the gaze vector estimated by a gaze estimation pipeline, is one of the most important ET signal quality characteristics.
-The subset of $8615$ participants who were recorded for both random saccade tasks were partitioned based on their average spatial accuracy during the two tasks.
+Although a constant error bias in the gaze estimates would not affect measures of velocity, we consider that spatial accuracy is often dependent on gaze angle and might therefore affect features such as saccade curvature.
+The subset of $8615$ participants from the full dataset who were recorded for both random saccade tasks were partitioned based on their average spatial accuracy during the two tasks.
 These partitions (illustrated in Figure 7) formed high error and low error train sets each with $3940$ participants, as well as high error, medium error, and low error test sets each with $245$ participants.
-Figure 7: Partitions of users in GazePro based on their average spatial accuracy.
-The degree value above a partition represents the median spatial accuracy across users in that partition.
-Note that only the subset of $8615$ users who were recorded for both random saccade tasks are included in these partitions.
+Figure 7: Partitions of a subset of users in GazePro based on their average spatial accuracy. The degree value above a partition represents the median spatial accuracy across users in that partition.
 EER was highest on the high error test set for both Experiments 16 and 17, which may indicate that poor spatial accuracy reduces authentication performance.
 Additionally, a model trained on data with high/low error performed worse when evaluated on data with low/high error.
+However, we observed a different trend when using the visual axis alone instead of both axes (see the supplementary material).
 (RQ7–8) Enrollment/verification duration.
-The experiments described above employed $20$ seconds of data for both enrollment and verification.
-However, it is important to understand how authentication accuracy changes for shorter presentation durations.
+The prior experiments employed $20$ seconds of data for both enrollment and verification.
 In this section, we employed presentation durations of $5$, $10$, and $15$ seconds to establish corresponding performance baselines.
-To do this, instead of computing a centroid embedding over the first four $5$-second chunks of a recording, we used the first one, two, or three chunks.
+To do this, instead of aggregating embeddings over the first four $5$-second chunks of a recording, we used the first one, two, or three chunks.
 We used the model from Experiment 5 for this analysis.
@@
 Using the first $15$ seconds for enrollment and verification, we achieved $0.05$% EER and $8.8$% $\text{FRR}_{0.002\%}$.
-If we instead used a fixed enrollment duration of $30$ seconds and only varied the verification duration, we observed a similar trend of authentication performance improving with increasing verification duration.
+If we instead used a fixed enrollment duration of $30$ seconds, we observed a similar trend of authentication performance improving with increasing verification duration.
 We observed that enrolling with $30$ seconds and verifying with $X$ seconds (up to $X=15$) yielded roughly similar results as enrolling and verifying with $X+5$ seconds.
 (RQ9) Task-independence.
-The prior experiments used a random saccade (jumping dot) task for both enrollment and verification.
+The prior experiments used a random saccade (jumping dot) task for enrollment and verification.
 How much, if at all, would performance change by using a smooth pursuit (gliding dot) task instead?
@@
 We found that the median ICC of the $128$ embedding features was $0.94$, with a minimum of $0.90$ and a maximum of $0.98$.
-Even our lowest ICC of $0.90$ meets all but the strictest of the commonly used thresholds [42], so all $128$ embedding features may be considered highly reliable.
 5 Discussion
@@
 The difference between the visual and optical axes would essentially provide an estimate of a user’s calibration parameters, and this difference alone achieves good verification accuracy (Experiment 4).
+These results suggest that the angle kappa between the optical and visual axes might be a highly distinctive trait.
 In a scenario such as continuous authentication where a calibration period could not precede every verification attempt, it may be feasible to employ a cached set of calibration parameters from the enrolled user to estimate the visual axis.
@@
 (RQ4) We observed an increase in verification rates as the training population size increased from $250$ to $6000$ users (Experiments 7–14), with the largest performance improvement occurring between $2000$ and $3000$ users.
-These results suggest that, if there is a point at which performance stops improving when more users are included in the training set, such a point would be beyond $6000$ users.
-Also, it is possible that training on an even larger population may yield even better verification rates.
+These results suggest that, if there is a point at which performance stops improving when more users are included in the training set, such a point would be beyond $6000$ users, and training on an even larger population may yield even better verification rates.
 (RQ5) We empirically confirmed in practice that, as the test population size increases, verification rates remain stable whereas identification rates decrease (Experiment 15).
 Eye movement traits are behavioral rather than physical, so they are expected to lack the distinctiveness necessary for large-scale identification.
-By fitting a curve of the form $a\sqrt{x}+b$ to the observed decline of identification rates and assuming the trend continues, we might expect random-chance identification rates at gallery sizes of about $148000$ identities (see Figure 6).
-(RQ6) Regarding the impact of spatial accuracy on authentication performance, we found that poor spatial accuracy reduced authentication performance (Experiments 16–17), in contrast to a study by Raju et al.
+(RQ6) Regarding the impact of spatial accuracy on authentication performance, we found that in some cases poor spatial accuracy reduced authentication performance (Experiments 16–17), in contrast to a study by Raju et al.
 [33].
 Our results also suggest that, for the best performance, the spatial accuracy of the training dataset should be representative of the level of spatial accuracy present in the testing dataset.
+Further research into the relationship between authentication performance and spatial accuracy is necessary.
 (RQ7–8) We observed that, using the model from Experiment 5, $\text{FRR}_{0.002\%}$ is approximately halved for every additional $5$ seconds used for enrollment and verification, starting at $41.9$% with $5$ seconds and improving to $4.1$% with $20$ seconds.
-Additionally, increasing the enrollment duration improves performance for shorter verification durations.
+Additionally, we observed that increasing the enrollment duration improves performance for shorter verification durations, potentially reducing the time needed for verification attempts at the cost of a longer initial enrollment procedure.
 It is important to note, though, that user calibration parameters have still been determined from a full-length calibration process.
@@
 This is evidence that gaze-based verification remains to be task-independent even for larger population sizes.
-(RQ10) The learned embedding features of the model from Experiment 5 all have ICCs at or above $0.9$, indicating that all $128$ embedding features are highly reliable.
+(RQ10) All 128 learned embedding features of the model from Experiment 5 have ICCs at or above $0.9$, indicating that they are highly reliable according to all but the strictest of the commonly used thresholds [42].
+This supports the idea from Friedman et al.
+[10] that higher ICCs are correlated with higher authentication accuracies.
 Further discussion of the limitations of this work can be found in the supplementary material.
@@
 In addition, we reinforced the importance of data length on resulting accuracy, specifically that longer enrollment/verification durations improved authentication accuracy.
-We also found that gaze-based authentication in our study appeared to be stimulus- and task-independent.
-In summary, gaze authentication can be as accurate as required by the FIDO standard when driven by a state-of-the-art machine learning architecture and a sufficiently large training dataset.
+We also found that gaze-based authentication in our study appeared to be task-independent.
+In summary, gaze authentication can be as accurate as required by the FIDO standard when driven by a state-of-the-art deep learning architecture and a sufficiently large training dataset.
 Acknowledgments
-We thank Meta for providing us with access to the GazePro dataset and compute resources to perform this important work.
+We thank Meta for providing us with access to the GazePro dataset and the compute resources necessary to perform this important work.